OpenStudy (anonymous):

Please explain how you find the answer for 5t^3 when t=2.

OpenStudy (anonymous):

Multiply t by itself 3 times ie 2*2*2 = 8 Now multiply this result by 5 ... 5 * 8 = 40

OpenStudy (amistre64):

t^3 means t*t*t 5*t*t*t ; when t=2 5*2*2*2 10*4 40

OpenStudy (amistre64):

does that make sense to you?

OpenStudy (anonymous):

Yes, it does. Thanks! I asked the wrong question however. What I meant to ask was, can you explain how to get the answer for 5(2+h)^3